Hi, Here are the questions written by my child's teacher- and we couldn't find any examples in the book just like these- please tell me if we did them correctly. I really don't like it when teachers start making up their own problems if there are not examples in the book for me to help my child with! Thank you!

Simplify, Show your work (what work?!)
(6)/(-2)= -3
(6)(2) = +12
(-6)/(-2)= +3
(6)(2)= -12

Answer by tutorcecilia(2152)

(6)/(-2)= (2)(3)/(-1)(2)= 3/-1= -3
.
(6)(2) = +12 (OK)
.
(-6)/(-2)= (-1)(2)(3)/(-1)(2) = 3
.
(6)(2)= 12 (not -12)
.
Example: (-6)(-12) = 12
Example: (-6)(12) = -12
Please let me know if you need any additional assistance. Thanks.